Find the critical numbers of the function. (Enter your answers as a comma-separated list. If an answer does not exist, enter DNE.)

f(x) = x3 + 9x2 − 81x
Answer #1

f(x) = x3 + q x² 81x f69 = 3*+10-101 1 f(X) = 3x2+10x-81 7 for critical numbers Junction 3x²+ 10x dl=0 x ²4 6x-28 =0 X?+ 9xelthen X+9 = 6 [x=-91 X-3 20 [x=3) Calculating t o filan = 6*418 at x = -9 f(x) = -54+18=-36 <D Therefore maxima at x = -9 at11 maxinuum nalle y fx at tangan f697 = €993 + q (+972-8 1-9 E-729 +229 +729 = 729 minimum values of for at x =43 163) = 3934